OpenStudy (anonymous):

Which value for the number makes this statement true? The quotient of a number and five is eight. 4 8 40 80

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

mathematically your statement should be written as: number/5 equals to 8 ..... n/5=8 multiply both sides by 5 n/5*5=8*5 n=40

OpenStudy (anonymous):

40 is my guess...

OpenStudy (anonymous):

you are right @johnnyy999

OpenStudy (anonymous):

Sweet thanks!

OpenStudy (anonymous):

np :)

TheSmartOne (thesmartone):

ksanka was able to help you :)